NEW BAND SYSTEMS OF YbI MOLECULE IN THERMAL EMISSION

\underline{K.N. UTTAM}, PAVITRA TANDON and RENU SINGH, Saha’s Spectroscopy Laboratory, Department of Physics, University of Allahabad, Allahabad (U.P.), India.

Thermal emission spectrum of Ytterbium monoiodide (YbI) molecule has been photographed for the first time in the spectral region 4500-5400 Å at a reciprocal linear dispersion of 7.3 Å/mm using high temperature vacuum graphite furnace at 2000 ^{∘}C and 2m Plane Grating Spectrograph (PGS-2). About 94 single and double headed, violet degraded bands have been recorded out of which 82 are entirely quite new. The observed bands have been classified into three systems out of which two are entirely new. The vibrational analysis has been carried out and it is suggested that these systems arise from the ground state ^{2}Σ. The vibrational constants determined for the systems are given below:

System  Region  ω′_{e}  ω′_{e}x′_{e} ω′′_{e} ω′′_{e}x′′_{e}  ν_{00}
B_{1}-X λλ5000 - 5300 Å 168.0 0.60 153.0 0.70 19133.6
B_{2}-X λλ5000 - 5200 Å 165.0 0.30 153.0 1.40 19381.0
C_{1}-X λλ4900 - 5100 Å 174.5 0.10 153.0 0.70 19940.2